OpenStudy (kylav210):

What is the magnitude and direction of Vector GH if G(2, -2) and H(7, 6)? magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (amistre64):

first off, lets position this at the origin by subtracting one point from the other

OpenStudy (amistre64):

H(7, 6) -G(2, -2) --------- (5, 10) this is the position representation of the vector from G to H

OpenStudy (amistre64):

since 5,10 is greater than 10,10 (45 degrees) its safe to say that the direction part is 55 or so 

OpenStudy (kylav210):

Okay

OpenStudy (amistre64):

the magnitude is just the length of the vector using the distance formula

OpenStudy (amistre64):

square add sqrt .... (5, 10) (5, 10)^2 = 25,100 25+100 = 125 sqrt(125) HMMM

OpenStudy (amistre64):

caps got stuck :/

OpenStudy (kylav210):

Thats okay , no problem!(:

OpenStudy (amistre64):

it appears that the magnitude doesnt fit any of the option .... sqrt(125) is about 11 or so

OpenStudy (amistre64):

pfft, 6+2 = 8 .. not 10 :) forgot how to add

OpenStudy (kylav210):

Lol!!

OpenStudy (amistre64):

so id still go with the larger 9 spot and about 55 for the win

OpenStudy (kylav210):

Okay! Thanks(:

OpenStudy (amistre64):

youre welcome, and with any luck, some of that made sense ;)

OpenStudy (kylav210):

Yes it did ! Thank you(: